Heather's school is charging $12 per uniform plus $200 fees and Brandon's School is charging $14 per uniform plus $196 fees
C = 12n + 200
C = 14n + 196
When will the two equations break even? Use substitution or elimination method

Answer by ewatrrr(24785)

 
Hi
C = 12n + 200
C = 14n + 196
C = 12n + 200
-C = -14n - 196
0 = -2n - 4
n = 2
If Heather and Brandon each purchased 2 uniforms, their cost would be the same
12*2 + 200 = $224 = 14*2 + 196
Heather Brandon